1.)Use De Moivre’s Theorem to

a.)derive the 4th roots of w =-8i

b.) express cos(4θ) and sin(5θ) in terms of powers of cos θ and sin θ

c.) expand cos^6 θ in terms of multiple powers of z based on θ

d.) express cos3θ sin4 θ in terms of multiple angles.

(2.)
a.) Let z =z1/z2 where z1 = tan θ + i and z2 = z1. Find an expression for z^n with n ∈ N.
b.) Let z = cos θ - i(1 + sin θ). Determine

2z + i/-1 - iz
1
Expert's answer
2021-07-22T06:38:02-0400

Solution.

1a

Derive the 4th roots of W=-8i

Write the complex number in polar form

If W=a+bi=rcos"\theta" +irsin"\theta"

r="\sqrt{a^2+b^2}=\sqrt{0+(-8)^2}=8"

"\theta=tan^-1(\frac{-8 }0)=-\fracπ2=-90^°"

W=8cos(-90°)+i8sin(-90°)

By De Moivres theorem,

W1/n=r1/n[cos"\frac{(2πk+\theta)}{n}+isin\frac{(2πk+\theta)}{n}]"

Where k=0,1,2...,n-1

1st root ,take k=0

"8^{\frac 1 4}[cos(-22.5°)+isin(-22.5°)]=1.554-0.6436i"

2nd root ,take k=1

"8^\frac1 4" [Cos(-22.5°+90°)+iSin(-22.5°+90°)]=0.6436+1.554i

3rd root ,take k=2

"8^\frac14"[Cos(-22.5+180°)+iSin (-22.5°+180°)]=-1.554+0.6436i

4th root,take k=3

8"^\frac14" [Cos(-22.5+270°)+iSin (-22.5°+270°)=-0.6436-1.554i


1b

De Moivres theorem;

"Cosn\theta+iSin n\theta=(cos\theta+isin\theta)^n"

Cos "4\theta=Re(cos\theta+iSin\theta)^4)"

"Cos4\theta=Re(Cos^4\theta+4Cos^3\theta(isin\theta)+6cos^2\theta(isin\theta)^2+4cos\theta(isin\theta)^3+(isin\theta)^4)"

"Cos4\theta=cos^4\theta-6cos^2\theta sin^2\theta+sin^4\theta"

sin"5\theta"

By De Moivre's Theorem;

"Cos5\theta+isin5\theta=(cos\theta+isin\theta)^5"

"Cos5\theta" +"isin5\theta" can now be expressed as;

"Cos5\theta+isin5\theta=cos^5\theta+5cos^4\theta isin\theta+10cos^3\theta i^2sin^2\theta+10cos^2\theta i^3sin^3\theta+5cos\theta i^4sin^4\theta+i^5sin^5\theta"

Group real parts and imaginary parts together;

"Cos5\theta+isin5\theta=(cos^5\theta-10cos^3\theta sin^2\theta+5cos\theta sin^4\theta)+i (5cos^4\theta sin\theta-10cos^2\theta sin^3\theta+sin^5\theta)"

Equate the imaginary parts of both sides together;

"isin5\theta=i(5cos^4\theta sin\theta-10cos^2\theta sin^3\theta+sin^5\theta)"

Divide out i

"Sin5\theta=5cos^4\theta sin\theta-10cos^2\theta sin^3\theta+sin^5\theta)"


1c

By De Moivre's Theorem;

zn+z-n=2cosn"\theta"

Take n=1

2cos"\theta" =z1+z-1

(2cos"\theta" )6=(z1+z-1)6

(2cos"\theta")6 =z6+6z4+15z2+20+15z-2+6z-4+z-6

64cos6"\theta" =(z6+z-6)+6(z4+z-4)+15(z2+z-2)+20

"64cos^6\theta=2cos6\theta+12cos4\theta+30cos2\theta+20"

Divide by 64;

"Cos^6\theta=\frac 2{32}cos6\theta+\frac3{16}cos4\theta+\frac{15}{32}cos2\theta+\frac5{16}"


1d

Find "Cos^3\theta"

(2cos"\theta")3=(z1+z-1)3=z3+3z1+3z-1+z-3

8cos3"\theta" =(z3+z-3)+3(z1+z-1)

=2"cos3\theta+6cos\theta"

cos"^3\theta=\frac14cos3\theta+\frac34cos\theta"

Find sin4"\theta"

(2isin"\theta" )4=(z-z-1)4=z4-4z2+6-4z-2+z-4

16sin"^4\theta" ="2cos4\theta-6cos2\theta+6"

"Sin^4\theta=\frac18cos4\theta-\frac12cos2\theta+\frac38"

Therefore;

"cos^3\theta sin^4\theta="("\frac14cos3\theta+\frac38cos\theta)(\frac18cos4\theta-\frac12cos2\theta+\frac38)"

Using the identity ;

CosAcosB="\frac12" [Cos(A+B)+Cos(A-B)

Simplify.

"\frac1{64}(cos7\theta-cos5\theta-3cos3\theta+3cos\theta)"


2a

"z =\frac{z_1}{z_2} \\
z_1 = tan θ + i \\
z_2 = \bar{z_1} = \bar{(\tan\theta+i)}=tan \theta -i\\
z =\frac{z_1}{z_2} =\frac{tan \theta +i}{tan \theta -i}\\
z =\frac{tan \theta +i}{tan \theta -i}*\frac{tan \theta +i}{tan \theta +i}\\
Simplifying \\
z=- cos 2\theta + sin2 \theta i\\
Hence\\
e^{i \theta}= cos \theta + isin \theta\\
e^{-i \theta}= cos \theta - isin \theta\\
\implies e^{i2\theta}= cos 2\theta + isin 2\theta\\
\space \space \space \space \space \space \space \space \space \space e^{-i \theta}= cos 2\theta -isin2 \theta\\
Adding \space them\\
cos 2 \theta = \frac{e^{i2 \theta }-e^{-i2 \theta }}{2}\\
Also, sin 2 \theta = \frac{e^{i2 \theta }-e^{-i2 \theta }}{2i}\\
z= \frac{e^{i2 \theta }-e^{-i2 \theta }}{2}+ \frac{e^{i2 \theta }-e^{-i2 \theta }}{2i}\\
z= -e^{i2 \theta }\\
\therefore z^n= (-e^{i2 \theta })^n = (-)^ne^{i2n \theta }"


2b

"z = cos θ - i(1 + sin θ)\\
|\frac{\bar{2z+i}}{-1-iz}|=|\frac{\bar{2(cos θ - i(1 + sin θ))+i}}{-1-i(cos θ - i(1 + sin θ))}|\\
=\frac{\bar{2cos θ - i(1 + 2sin θ)}}{(-2- sin θ)-icos \theta}*\frac{-2-sin \theta+icos \theta}{(-2- sin θ)-icos \theta}\\
= \frac{-3 cos \theta + i (5 sin \theta +3)}{4 sin \theta +5}\\
Conjugate \\
\frac{-3 cos \theta + i (5 sin \theta +3)}{4 sin \theta +5}\\
|\frac{\bar{2z+i}}{-1-iz}|= \sqrt{(\frac{-3 cos \theta }{4 sin \theta +5})^2+(\frac{5 sin \theta +3 }{4 sin \theta +5})^2}\\
= \frac{ \sqrt{16 sin^2+30sin \theta +18}}{4 sin \theta +5}"